OpenStudy (anonymous):

lim as x approaches negative infinity of x+sqrt(x^2+2x)

OpenStudy (anonymous):

So I multiplied by the conjugate and got \[\frac{ 2x }{ x-\sqrt{x^2+2x} }\] not sure how to proceed from there.

OpenStudy (anonymous):

is that algebra right? let me check

OpenStudy (anonymous):

I think so.

OpenStudy (anonymous):

it maybe it is supposed to be -2x on top?

geerky42 (geerky42):

yeah, supposed to be -2x.

OpenStudy (anonymous):

still unsure what to do from there though

OpenStudy (anonymous):

no you are right

OpenStudy (anonymous):

divide by \(x\) i think

OpenStudy (anonymous):

oh yeah top should be \(-2x\) right now i think you can do it if you divide by \(x\)

OpenStudy (anonymous):

ok so then \[\frac{ -2 }{ 1-\sqrt{x^2+2x}/x }\]

OpenStudy (anonymous):

now if i put infinity in i get 1 - infinity on bottom so it is -2 over -infinity so negatives cancel and left with 2 over infinity so limit is 0?

OpenStudy (anonymous):

no hold on let me try this algebra, i think the denominator is going to be \(2\)

OpenStudy (anonymous):

can u show me how please

OpenStudy (anonymous):

yeah lets just look at the denominator

OpenStudy (anonymous):

you have \[\lim_{x\to -\infty}1+{\frac{\sqrt{x^2+2x}}{x}}\]

OpenStudy (anonymous):

nope i wrote it wrong

OpenStudy (anonymous):

\[\lim_{x\to -\infty}1-{\frac{\sqrt{x^2+2x}}{x}}\]

OpenStudy (anonymous):

now here is the thing, when you bring the \(x\) inside the radical you have to change the sign, since x is going to minus infinity not plus infinity

OpenStudy (anonymous):

but its squared

OpenStudy (anonymous):

yeah, the numerator is positive and the denominator is negative (since x is going to minus infinity)

OpenStudy (anonymous):

i mean of this portion \[\frac{\sqrt{x^2+2x}}{x}\]

OpenStudy (anonymous):

ok so how do you know which x is dominant term to plug infinity into

OpenStudy (anonymous):

or negative infintyy

OpenStudy (anonymous):

if you want to be very precise, you change the sign and bring the x in the denominator as \(x^2\) inside the radical and make it \[\lim_{x\to \infty}\sqrt{1+\frac{2}{x}}\]

OpenStudy (anonymous):

which is pretty clearly \(1\) making the whole denominator \(2\) and your final answer \(-1\)

OpenStudy (anonymous):

oh ok that makes sense

OpenStudy (anonymous):

hope so good luck!